A rancher is planning to put up 220 yards of fencing. In the morning she puts up 80 yards, and in the afternoon she puts up 40% of the remaining fence.What percent of the fence did she put up that day?
Answer by xcentaur(357)

total=220 yards
morning done=80 yards
remaining=220-80=140 yards
afternoon done=40% of remaining
=40/100*140
=56 yards
total done=morning+afternoon=80+56=136 yards


Percentage done=136/220*100
=61.81%


Hope this helps,
good luck